Please help me

substitution
x=y+7
y-8=2x

Answer by jim_thompson5910(35256)

Replace EVERY x in y-8=2x with y+7. Why is this valid? Because we're given that x = y+7

So we go from

y-8=2x

to

y-8=2(y+7)


From here, we distribute

y-8 = 2y + 14

I'll let you finish up. You need to solve for y. Once you have the solution for y, you plug it into x = y+7 to get the solution for x.